OpenStudy (anonymous):

Can you help me solve for x? x^2-8x+15=0

OpenStudy (anonymous):

\[\huge x^2-8x+15=>x^2-5x-3x+15=0\]

OpenStudy (anonymous):

\[\huge x(x-5)-3(x-5)=(x-3)(x-5)\]

OpenStudy (anonymous):

\[x=3&5\]